Circular No. 8168

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2003gn, 2003go, 2003gp, 2003gq Four additional apparent supernovae have been found on unfiltered

CCD images, each by LOTOSS/KAIT (via J. Graham, D. Weisz, and W. Li; cf. IAUC 8162), with SN 2003gq reported initially by T. Puckett and A. Langoussis (cf. IAUC 8154; Puckett’s data for SN 2003gq tabulated below). Li provides position end figures 20^{s}.74, 57′′.6 for SN 2003gq.

SN  2003 UT  α_{2000}  δ_{2000}  Mag.  Offset
2003gn July 22.4 22^{h}33^{m}51.^{s}68 +20^{o}48′08′′.5 18.4 16′′.4 E, 11′′.0 S
2003go July 22.3 19 52 47.68 −19 11 28.2 18.6 4′′.8 E, 0′′.7 S
2003gp July 23.2 16 03 55.47 +25 00 24.8 18.6 12′′.9 E, 12′′.0 S
2003gq July 24.32 22 53 20.68 +32 07 57.6 17.6 5′′.0 W, 11′′.3 N

Additional approximate magnitudes, from KAIT unless otherwise noted: SN 2003gn, July 10.4 UT, [19.0; 17.4, [18.0; 23.4, 18.1. SN 2003go, July 17.3, [19.0; 23.3, 18.6. 2003gp in UGC 10160, July 16.2, [19.0; 25.2, 18.6. SN 2003gq in NGC 7407, July 8, [20.0 (Puckett); 16.4, 19.0; 25.22, 17.7 (M. Marcus and Puckett); 25.4, 17.6; 26.4, 17.5.

XTE J1810–197 C. B. Markwardt, University of Maryland and Goddard Space Flight

Center (GSFC); A. I. Ibrahim, GSFC and George Washington University; and J. H. Swank, GSFC, report that, following reports from the IPN and HETE of renewed bursting activity by the soft γ-ray repeater SGR 1806– 20 (GCN 2297, 2308), an RXTE PCA pointed observation was undertaken that revealed unexpected x-ray pulsations at a period of 5.54 s in the PCA field-of-view (GCN 2306). An RXTE PCA scanning observation on July 18.23–18.27 UT showed a new source at α = 18^{h}10^{m}.9, δ = −19^{o}42′ (equinox 2000.0), with estimated 99-percent-confidence semi-major axes of 5′ in α and 7′ in δ; this confidence region excludes both the reported IPN burst annuli and the HETE burst error region, and is not consistent with any other known pulsar. The pulsations are clearly from this source, designated XTE J1810–197. The fitted flux was 2.5 ± 0.2 mCrab (2–10 keV). On hindsight, XTE J1810–197 has been detectable in RXTE PCA galacticbulge scans since Feb. 7 (but not before), and it is steadily declining from a maximum of 3.6 mCrab. Follow-up observations are encouraged.

2003 July 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green